# Skullptor: High Fidelity 3D Head Reconstruction in Seconds with Multi-View Normal Prediction

### v1

@@ Abstract
Reconstructing high-fidelity 3D head geometry from images is critical for a wide range of applications, yet existing methods face fundamental limitations. Traditional photogrammetry achieves exceptional detail but requires extensive camera arrays (25-200+ views), substantial computation, and manual cleanup in challenging areas like facial hair. Recent alternatives present a fundamental trade-off: foundation models enable efficient single-image reconstruction but lack fine geometric detail, while optimization-based methods achieve higher fidelity but require dense views and expensive computation. We bridge this gap with a hybrid approach that combines the strengths of both paradigms. Our method introduces a multi-view surface normal prediction model that extends monocular foundation models with cross-view attention to produce geometrically consistent normals in a feed-forward pass. We then leverage these predictions as strong geometric priors within an inverse rendering optimization framework to recover high-frequency surface details. Our approach outperforms state-of-the-art single-image and multi-view methods, achieving high-fidelity reconstruction on par with dense-view photogrammetry while reducing camera requirements and computational cost. The code and model will be released.
Figure 1: Given sparse multi-view images of a subject, our method predicts geometrically consistent surface normals and leverages them to recover complete and detailed 3D head geometry preserving fine surface features such as wrinkles and skin folds in a matter of seconds.
@@ 1 Introduction
Photogrammetry [26, 61, 19] has long served as the gold standard for acquiring detailed 360 degree head geometry in production pipelines, making it the preferred choice for high-end VFX and gaming applications. However, it requires dense camera coverage (typically 50-200+ synchronized views), extensive processing time, and substantial computational resources. For 4D sequences, the storage requirements become prohibitive with multi-view capture, generating terabytes of data. Moreover, photogrammetry reconstruction often struggles with the presence of view-dependent appearance (e.g., specular reflections) and fine geometry such as facial hair, leading to artifacts and incomplete geometry that require manual correction.
Recently, two technical approaches have emerged to reduce the processing time and the number of viewpoints required for high-fidelity 3D reconstruction;however, both navigate a fundamental trade-off between speed, ease of data capture, and geometric accuracy. The first leverages data-driven foundation models for human vision [36, 55, 48, 65], which recover 3D geometry from a single image in an efficient feed-forward pass, dramatically simplifying data capture compared to photogrammetry. Still, even state-of-the-art foundation models [36] fall short of photogrammetry in reconstructing fine geometric detail (Figure 3), likely because they rely on learned, and often ambiguous, 3D shape priors rather precise multi-view geometric constraints.
A second approach reconstructs 3D geometry through direct, iterative optimization of implicit or explicit 3D representations [24, 30, 54]. These representations are easier to optimize than traditional mesh-based formulations and achieve higher geometric fidelity than foundation models by explicitly enforcing multi-view consistency, yet they still struggle to capture high-frequency geometric details such as wrinkles, skin folds, and person-specific surface variations (Figure 4). Moreover, because they lack strong data-driven priors, they typically require extensive camera arrays to provide dense view coverage, as well as computationally expensive optimization procedures. Consequently, this class of methods lags behind foundation models in terms of computational efficiency and ease of capture. Overall, neither photogrammetry nor these emerging methods achieve all three desiderata simultaneously: high geometric accuracy, a sparse view capture setup, and computationally efficient reconstruction.
@@
  • We develop a multi-view normal prediction model that efficiently adapts a monocular facial foundation model (DAViD [55]) using a lightweight view-aware cross-attention mechanism to enforce geometric consistency, generating a significantly more accurate surface normals from sparse (<<10) input viewpoints.
@@ 2 Related Work
Multi-view stereo (photogrammetry) [26, 61, 19], implemented in open frameworks like Colmap [56, 57] and Meshroom [23], as well as commercial solutions,111 RealityCapture: https://www.capturingreality.com/ Agisoft: https://www.agisoft.com/, remains the gold standard for 3D facial capture in production. These methods achieve exceptional accuracy by triangulating 3D structure from dense correspondences across multiple views. However, photogrammetry requires dense camera coverage (25250+25-250+ views), substantial computational resources, and struggles with non-Lambertian surfaces and fine structures like facial hair. Recent approaches aim to simplify multi-view reconstruction through learned priors or flexible 3D representations, yet navigate a fundamental trade-off between geometric accuracy, computational efficiency, and capture requirements.
Data-driven methods Data-driven approaches prioritize computational efficiency by leveraging learned priors from large-scale datasets, typically trained on monocular 2D images. Although fast, these methods typically sacrifice geometric accuracy due to the inherent ambiguity in mapping 2D observations to 3D structure without explicit multi-view constraints.
@@
Monocular normal and depth estimation models offer greater flexibility than parametric models by predicting per-pixel geometric information without topological constraints. Recent estimators achieve strong performance through pre-training on domain-specific synthetic data [55, 36], leveraging large-scale diverse datasets [67, 68], or by fine-tuning general-purpose models [69] like Marigold [34] (adapted from Stable Diffusion) and Pixel3DMM [22] (built on DINOv2 [48]). DAViD [55] uses Dense Prediction Transformers [51], combining Vision Transformers [16] with convolutional blocks for high-resolution feature extraction. However, monocular estimators suffer from inherent single-image ambiguities and process each view independently, so geometric consistency across viewpoints is not guaranteed, leading to contradictory predictions that cannot be directly integrated into coherent 3D mesh reconstruction.
@@
Optimization-based methods Direct optimization methods reconstruct 3D geometry through iterative refinement of explicit or implicit representations. Early approaches use voxel grids [31] or meshes [10, 33, 63], but are limited by fixed topologies and memory constraints. Recent methods leverage flexible implicit representations such as signed distance functions [54, 53, 25] or combine reflectance and normal maps within volume rendering frameworks [8, 7]. Others explore explicit point-based representations, particularly 3D Gaussian Splatting [35], with extensions like SuGaR [24] and 2DGS [29] that extract surface meshes. [50] uses a prior head model from thousands of scans to overfit a neural network via differentiable rendering [47]. These methods achieve higher geometric fidelity than data-driven approaches by explicitly enforcing multi-view photometric consistency. However, without strong learned priors, they typically require dense view coverage and computationally expensive per-scene optimization, making them impractical for sparse-view scenarios.
Inverse rendering Differentiable rendering libraries like such [52, 38] enable inverse rendering by optimizing mesh geometry and appearance to match observed images. This approach has a long history, with foundational work establishing energy functions that balance data fidelity with geometric regularization [27, 5, 60, 32]. However, directly applying gradient updates with standard optimizers like ADAM [37] often leads to mesh degradation, including self-intersections and flipped triangles. While periodic remeshing [6, 2] can mitigate these issues, it disrupts the optimization process. Advanced techniques such as Continuous Remeshing [49] integrate adaptive topology management directly into each optimization iteration, preventing mesh defects and enabling stable refinement of high-frequency geometric details.
Our approach
@@ 3.1 Consistent Multi-View Normal Prediction
Training The model takes $m$ input images $\mathcal{I}$ and predicts their corresponding normal maps $\hat{\mathcal{N}}$ at the same spatial resolution. We use cosine similarity loss as the objective function:
@@
where $\hat{\mathcal{N}}_{i}$ is the predicted normal or the $i$th view and $\mathcal{N}_{i}$ is the corresponding ground truth normal.
@@
Rather than training on images from fixed lightstage camera positions, we render the ground truth geometry (with its corresponding texture) from synthesized viewpoints with randomly sampled virtual cameras. For each training sample, we render both the textured image $\mathcal{I}_{i}$ and its corresponding ground truth normal map $\mathcal{N}_{i}$ (in camera space) from the head geometry. This rendering-based data augmentation provides two key advantages: (1) it exposes the model to diverse camera configurations beyond the lightstage setup, improving generalization to arbitrary multi-view capture systems, and (2) it increases training data diversity without requiring additional physical captures. As we demonstrate in section 4, our model is able to generalize effectively to different capture setups, including NPHM and Multiface datasets with different camera configurations. Technical details are provided in the supplementary material.
@@ 3.2 Normal-Guided Mesh Optimization
Given the predicted multi-view surface normals $\hat{\mathcal{N}}$ (subsection 3.1), we recover detailed 3D facial geometry through optimization-based mesh refinement. Our approach iteratively adjusts mesh vertex positions such that the rendered surface normals match the predicted normals under differentiable rendering. A key challenge is establishing geometric correspondence between the predicted normals, expressed in local camera coordinates, and a consistent global coordinate space.
@@
Mesh optimization We initialize the mesh $\mathcal{M}$, that we want to optimize, to a unit sphere, and optimize it to minimize the distance between the predicted normals $\hat{\mathcal{N}}^{*}$ (from previous stage) and the normals rendered $\bar{\mathcal{N}}$ from the $\mathcal{M}$. At each step of the optimization, we render the mesh under the transformed camera parameters using a differentiable rendering function $\mathcal{F}$ defined as follows:
@@
where $\hat{\mathcal{W}}_{i}$ is a per-pixel weight matrix for the view $i$. This weight matrix prioritizes regions facing the camera where predictions are more reliable, defined at a given pixel position $(u,v)$ by:
@@
At the end of each optimization step, we apply adaptive remeshing operations from  [49]. These operations consists of edge splits, collapses, and flips that dynamically adjust mesh resolution to local geometric complexity while preventing degeneracies such as collapsed faces and self-intersections. The final output is a normalized mesh aligned to the canonical coordinate system. While we initialize $\mathcal{M}$ with a unit sphere for simplicity, using a aligned template head mesh as initialization did not show meaningful differences in terms of reconstruction time or quality. More details on the optimization in the supp. material.
@@ 4 Evaluation
The evaluation of our proposed pipeline aims to validate three key claims: (1) our extension of monocular normal estimation models to multi-view processing improves normal map estimation, especially for capturing high-frequency details; (2) our optimization-based remeshing effectively leverages these multi-view normal estimates to produce a detailed mesh; and (3) the complete pipeline can outperform standard photogrammetry in quality when using sparse camera arrays (<10<10), while also being more computationally efficient.
@@ 4.1 Experimental Setup
Figure 3: Qualitative comparison of normal predictions across different methods. Top 2: NPHM; Bottom 2: Multiface
Datasets
@@
NPHM: This dataset contain highly detailed 3D head scans of 124 subjects, each performing 20 distinct facial expressions, captured with a custom structured-light scanning setup, using Artec Eva scanners [59]. This capture setup provides highly accurate geometry estimations without using a photogrammetry pipeline, making it ideal for evaluations involving photogrammetry. The dataset does not contain captured images, so we generate a synthetic dataset by rendering NPHM scans from 23 virtual camera viewpoints. We render normal maps and textured images to serve as ground truth and input data, respectively. For a tractable evaluation, we use a subset consisting of the first 20 subjects, for a total of 400 expression scans.
@@
Multiface This dataset is composed of lightstage multi-view video recordings of 8 subjects from 40 different cameras performing a variety of facial expressions, resulting in thousands of dynamic frames. For this dataset, we directly use the captured images as input. Ground-truth geometry is obtained via a high-quality photogrammetry reconstruction pipeline, producing dense, high-fidelity meshes for each frame. To match the scale of our NPHM dataset, we subsample the sequences by selecting 1 out of every 30 frames, yielding a diverse evaluation set of 347 frames in total.
@@
For both datasets, all ground-truth meshes are aligned to a canonical coordinate system using the same technique described in subsection 3.2. This ensures all reconstructions are in a canonical space, allowing for consistent results on metrics that depend on scale, such as depth error. Further details on how the evaluation dataset was created is available in the supp. material.
@@
Baselines: We compare our method against several state-of-the-art baselines at two different stages:
@@
  • Normal estimation: We compare our multi-view normal estimator against leading monocular methods, specifically, the foundation model Sapiens[36] using both its lightweight 0.3B architecture, as well as its full 2B parameter mode, and the original DAViD model [55].
@@
  • Full mesh reconstruction: We compare our complete pipeline against established multi-view reconstruction methods: Meshroom [23] (a robust open-source photogrammetry pipeline) and recent Gaussian splatting-based methods 2DGS [29] and SuGaR [24], which recover explicit geometry from gaussian splats.
@@
Evaluation Metrics. We evaluate normal prediction quality using: (1) Mean Angular Difference (in degrees), measuring the angular error between predicted and ground truth normals; (2) Percentage of angles below 10°, 20°, and 30°, measuring the fraction of predictions whose angular error falls within these thresholds; and (3) Normal Gradient Error, a novel metric that captures preservation of high-frequency geometric details. This metric applies a Sobel filter to both predicted and ground truth normals and measures the L1 distance between filtered results. This metric directly measures the preservation of fine geometric details by capturing local surface curvature variations. Since normals are spatial derivatives of depth, their gradients emphasize high-frequency features such as wrinkles, folds and subtle surface variations, that are critical for photorealistic reconstruction. Methods that produce overly smooth geometry may achieve low angular error while scoring poorly on normal gradient error, making this metric particularly relevant for evaluating our approach’s ability to capture fine-scale detail. Additionally, for mesh evaluations, we include a L1 depth error to assess the global quality of the mesh reconstruction.
4.2 Normal Evaluation
@@
* Out of the 50 subjects from each dataset, we excluded 6 additional subjects from Multiface and 12 from NPHM due to failed reconstructions from 2DGS and SuGaR, in order to provide a fair comparison.
In this section, we validate the effectiveness of our multi-view normal estimation network (Sec. 3.1). We compare it against the monocular baselines defined in subsection 4.1. As ground truth, we rasterize normals from the ground truth meshes with camera views identical to the ones used for input images. Additionally, we use the Facer face segmentation [70] to filter out hair, cloth and neck of each subject, ensuring that normal prediction metrics are calculated only on the face. We also report the average inference time for predicting 1 batch of 10 images.

### v3

@@ Abstract
Reconstructing high-fidelity 3D head geometry from images is critical for a wide range of applications, yet existing methods face fundamental limitations. Traditional photogrammetry achieves exceptional detail but requires extensive camera arrays (25–200+ views), substantial computation, and manual cleanup in challenging areas like facial hair. Recent alternatives present a fundamental trade-off: foundation models enable efficient single-image reconstruction but lack fine geometric detail, while optimization-based methods achieve higher fidelity but require dense views and expensive computation. We bridge this gap with a hybrid approach that combines the strengths of both paradigms. Our method introduces a multi-view surface normal prediction model that extends monocular foundation models with cross-view attention to produce geometrically consistent normals in a feed-forward pass. We then leverage these predictions as strong geometric priors within an inverse rendering optimization framework to recover high-frequency surface details. Our approach outperforms state-of-the-art single-image and multi-view methods, achieving high-fidelity reconstruction on par with dense-view photogrammetry while reducing camera requirements and computational cost. Results available on the project page.
Figure 1: Given sparse, multi-view images of a subject, our method predicts geometrically consistent surface normals and leverages them to recover complete and detailed 3D head geometry, recovering fine surface features such as wrinkles and skin folds in a matter of seconds.
@@ 1 Introduction
Photogrammetry [26, 60, 19] has long served as the gold standard for acquiring detailed 360-degree head geometry in production pipelines, making it the preferred choice for high-end VFX and gaming applications. However, it requires dense camera coverage (typically 50–200+ synchronized views), extensive processing time, and substantial computational resources. For 4D sequences, the storage requirements become prohibitive with multi-view capture, generating terabytes of data. Moreover, photogrammetry reconstruction often struggles with the presence of view-dependent appearance (e.g., specular reflections) and fine geometry such as facial hair, leading to artifacts and incomplete geometry that require manual correction.
Recently, two technical approaches have emerged to reduce the processing time and the number of viewpoints required for high-fidelity 3D reconstruction; however, both navigate a fundamental trade-off between speed, ease of data capture, and geometric accuracy. The first leverages data-driven foundation models for human vision [35, 54, 47, 64], which recover 3D geometry from a single image in an efficient feed-forward pass, dramatically simplifying data capture compared to photogrammetry. Still, even state-of-the-art foundation models [35] fall short of photogrammetry in reconstructing fine geometric detail, likely because they rely on learned, and often ambiguous, 3D shape priors rather than precise multi-view geometric constraints.
A second approach reconstructs 3D geometry through direct, iterative optimization of implicit or explicit 3D representations [24, 29, 53]. These representations are easier to optimize than traditional mesh-based formulations and achieve higher geometric fidelity than foundation models by explicitly enforcing multi-view consistency, yet they still struggle to capture high-frequency geometric details such as wrinkles, skin folds, and person-specific surface variations. Moreover, because they lack strong data-driven priors, they typically require extensive camera arrays to provide dense view coverage, as well as computationally expensive optimization procedures. Consequently, this class of methods lags behind foundation models in terms of computational efficiency and ease of capture. Overall, neither photogrammetry nor these emerging methods achieve all three desiderata simultaneously: high geometric accuracy, a sparse view capture setup, and computationally efficient reconstruction.
@@
  • We develop a multi-view normal prediction model that efficiently adapts a monocular facial foundation model (DAViD [54]) using a lightweight view-aware cross-attention mechanism to enforce geometric consistency, generating significantly more accurate surface normals from sparse (<<10) input viewpoints.
@@ 2 Related Work
Multi-view stereo (photogrammetry) [26, 60, 19], implemented in open frameworks like COLMAP [55, 56] and Meshroom [23], as well as commercial solutions111 RealityCapture: https://www.capturingreality.com Agisoft: https://www.agisoft.com , remains the gold standard for 3D facial capture in production. These methods achieve exceptional accuracy by triangulating 3D structure from dense correspondences across multiple views. However, photogrammetry requires dense camera coverage (25–250+ views), substantial computational resources, and struggles with non-Lambertian surfaces and fine structures like facial hair. Recent approaches aim to simplify multi-view reconstruction through learned priors or flexible 3D representations, yet navigate a fundamental trade-off between geometric accuracy, computational efficiency, and capture requirements.
Data-driven methods. Data-driven approaches prioritize computational efficiency by leveraging learned priors from large-scale datasets, typically trained on monocular 2D images. Although fast, these methods typically sacrifice geometric accuracy due to the inherent ambiguity in mapping 2D observations to 3D structure without explicit multi-view constraints.
@@
Monocular normal and depth estimation models offer greater flexibility than parametric models by predicting per-pixel geometric information without topological constraints. Recent estimators achieve strong performance through pre-training on domain-specific synthetic data [54, 35], leveraging large-scale diverse datasets [66, 67], or by fine-tuning general-purpose models [68] like Marigold [33] (adapted from Stable Diffusion) and Pixel3DMM [22] (built on DINOv2 [47]). DAViD [54] uses Dense Prediction Transformers [50], combining vision transformers [16] with convolutional blocks for high-resolution feature extraction. However, monocular estimators suffer from inherent single-image ambiguities and process each view independently, so geometric consistency across viewpoints is not guaranteed, leading to contradictory predictions that cannot be directly integrated into a coherent 3D mesh reconstruction.
@@
Optimization-based methods. Direct optimization methods reconstruct 3D geometry through iterative refinement of explicit or implicit representations. Early approaches use voxel grids [30] or meshes [10, 32, 62], but are limited by fixed topologies and memory constraints. Recent methods leverage flexible implicit representations such as signed distance functions [53, 52, 25] or combine reflectance and normal maps within volume rendering frameworks [8, 7]. Others explore explicit point-based representations, particularly 3D Gaussian Splatting [34], with extensions like SuGaR [24] and 2DGS [29] that extract surface meshes. Ramon et al. [49] use a prior head model from thousands of scans to overfit a neural network via differentiable rendering [46]. These methods achieve higher geometric fidelity than data-driven approaches by explicitly enforcing multi-view photometric consistency. However, without strong learned priors, they typically require dense view coverage and computationally expensive per-scene optimization, making them impractical for sparse-view scenarios.
Inverse rendering. Differentiable rendering libraries such as [51, 37] enable inverse rendering by optimizing mesh geometry and appearance to match observed images. This approach has a long history, with foundational work establishing energy functions that balance data fidelity with geometric regularization [27, 5, 59, 31]. However, directly applying gradient updates with standard optimizers like ADAM [36] often leads to mesh degradation, including self-intersections and flipped triangles. While periodic remeshing [6, 2] can mitigate these issues, it disrupts the optimization process. Advanced techniques such as Continuous Remeshing [48] integrate adaptive topology management directly into each optimization iteration, preventing mesh defects and enabling stable refinement of high-frequency geometric details.
Our approach.
@@ 3.1 Consistent Multi-View Normal Prediction
Training. The model takes $m$ input images $\mathcal{I}$ and predicts their corresponding normal maps $\hat{\mathcal{N}}$ at the same spatial resolution. We use cosine similarity loss as the objective function:
@@
where $\hat{\mathcal{N}}_{i}$ is the predicted normal for the $i^{\text{th}}$ view and $\mathcal{N}_{i}$ is the corresponding ground truth normal.
@@
Rather than training on images from fixed lightstage camera positions, we render the ground truth geometry (with its corresponding texture) from synthesized viewpoints with randomly sampled virtual cameras. For each training sample, we render both the textured image $\mathcal{I}_{i}$ and its corresponding ground truth normal map $\mathcal{N}_{i}$ (in camera space) from the head geometry. This rendering-based data augmentation provides two key advantages: (1) it exposes the model to diverse camera configurations beyond the lightstage setup, improving generalization to arbitrary multi-view capture systems, and (2) it increases training data diversity without requiring additional physical captures. As we demonstrate in Sec. 4, our model is able to generalize effectively to different capture setups, including NPHM and Multiface datasets with different camera configurations. Technical details are provided in the supplementary material.
@@ 3.2 Normal-Guided Mesh Optimization
Given the predicted multi-view surface normals $\hat{\mathcal{N}}$ (Sec. 3.1), we recover detailed 3D facial geometry through optimization-based mesh refinement. Our approach iteratively adjusts mesh vertex positions such that the rendered surface normals match the predicted normals under differentiable rendering. A key challenge is establishing geometric correspondence between the predicted normals, expressed in local camera coordinates, and a consistent global coordinate space.
@@
Mesh optimization. We initialize the mesh $\mathcal{M}$, that we want to optimize, to a unit sphere, and optimize it to minimize the distance between the predicted normals $\hat{\mathcal{N}}^{*}$ (from previous stage) and the normals rendered $\bar{\mathcal{N}}$ from the $\mathcal{M}$. At each step of the optimization, we render the mesh under the transformed camera parameters using a differentiable rendering function $\mathcal{F}$ defined as follows:
@@
where $\hat{\mathcal{W}}_{i}$ is a per-pixel weight matrix for the view ii. This weight matrix prioritizes regions facing the camera where predictions are more reliable, defined at a given pixel position $(u,v)$ by
@@
At the end of each optimization step, we apply adaptive remeshing operations from  [48]. These operations consist of edge splits, collapses, and flips that dynamically adjust mesh resolution to local geometric complexity while preventing degeneracies such as collapsed faces and self-intersections. The final output is a normalized mesh aligned to the canonical coordinate system. While we initialize $\mathcal{M}$ with a unit sphere for simplicity, using an aligned template head mesh as initialization did not show meaningful differences in terms of reconstruction time or quality. More details are available in the supplemental material.
@@ 4 Evaluation
The evaluation of our proposed pipeline aims to validate three key claims: (1) our extension of monocular normal estimation models to multi-view processing improves normal map estimation, especially for capturing high-frequency details; (2) our optimization-based remeshing effectively leverages these multi-view normal estimates to produce a detailed mesh; and (3) the complete pipeline outperforms standard photogrammetry in quality when both methods operate on sparse camera arrays (<10<10), while also being more computationally efficient.
@@ 4.1 Experimental Setup
Figure 3: Qualitative comparison of normal predictions on the NPHM dataset (rows 1–2) and Multiface dataset (rows 3–4).
Datasets.
@@
NPHM. This dataset contains highly detailed 3D head scans of 124 subjects, each performing 20 distinct facial expressions, captured with a custom structured-light scanning setup using Artec Eva scanners [58]. This capture setup provides highly accurate geometry estimations without using a photogrammetry pipeline, making it ideal for evaluations involving photogrammetry. The dataset does not contain captured images, so we generate a synthetic dataset by rendering textured NPHM scans from 23 virtual camera viewpoints. We render normal maps and textured images to serve as ground truth and input data, respectively. For a tractable evaluation, we use a subset consisting of the first 20 subjects, for a total of 400 expression scans.
@@
Multiface. This dataset is composed of lightstage multi-view video recordings of 8 subjects from 40 different cameras performing a variety of facial expressions, resulting in thousands of dynamic frames. For this dataset, we directly use the captured images as input. Ground-truth geometry is obtained via a high-quality photogrammetry reconstruction pipeline, producing dense, high-fidelity meshes for each frame. To match the scale of our NPHM dataset, we subsample the sequences by selecting 1 out of every 30 frames, yielding a diverse evaluation set of 347 frames in total.
@@
For both datasets, all ground-truth meshes are aligned to a canonical coordinate system using the same technique described in Sec. 3.2. This ensures all reconstructions are in a canonical space, allowing for consistent results on metrics that depend on scale, such as depth error. Further details on how the evaluation dataset was created are available in the supplemental material.
@@
Baselines. We compare our method against several state-of-the-art baselines at two different stages:
@@
  • Normal estimation. We compare our multi-view normal estimator against leading monocular methods, specifically, the foundation model Sapiens [35] using both its lightweight 0.3B architecture, as well as its full 2B parameter mode, and the original DAViD model [54].
@@
  • Full mesh reconstruction. We compare our complete pipeline against established multi-view reconstruction methods: Meshroom [23] (a robust open-source photogrammetry pipeline) and recent Gaussian splatting-based methods 2DGS [29] and SuGaR [24], which recover explicit geometry from Gaussian splats.
@@
Evaluation metrics. We evaluate normal prediction quality using: (1) Mean Angular Difference (in degrees), measuring the angular error between predicted and ground truth normals; (2) Percentage of angles below 10°, 20°, and 30°, measuring the fraction of predictions whose angular error falls within these thresholds; and (3) Normal Gradient Error, a novel metric that captures preservation of high-frequency geometric details. This metric applies a Sobel filter to both predicted and ground truth normals and measures the L1 distance between filtered results. This metric directly measures the preservation of fine geometric details by capturing local surface curvature variations. Since normals are spatial derivatives of depth, their gradients emphasize high-frequency features such as wrinkles, folds and subtle surface variations that are critical for photorealistic reconstruction. Methods that produce overly smooth geometry may achieve low angular error while scoring poorly on normal gradient error, making this metric particularly relevant for evaluating our approach’s ability to capture fine-scale detail. Additionally, for mesh evaluations, we include an L1 depth error to assess the global quality of the mesh reconstruction.
@@
* Metrics exclude 6 Multiface and 12 NPHM subjects from the initial 50 faces per dataset where 2DGS or SuGaR failed to produce a reconstruction.
4.2 Normal Evaluation
In this section, we validate the effectiveness of our multi-view normal estimation network (Sec. 3.1). We compare it against the monocular baselines defined in Sec. 4.1. For ground truth, we rasterize normals from the ground truth meshes with camera views identical to the ones used for input images. Additionally, we use the Facer face segmentation [69] to filter out hair, cloth and neck of each subject, ensuring that normal prediction metrics are calculated only on the face. We also report the average inference time for predicting one batch of 10 images.